One hour after Yolanda started walking from X to Y, a distance of 45 miles, Bob started walking along the same road from Y to X. If Yolanda’s walking rate was 3 miles per hour and Bob’s was 4 miles per hour, how many miles had Bob walked when they met?
(A) 24
(B) 23
(C) 22
(D) 21
(E) 19.5

Answer by khwang(438)

You may have typing error about 45 miles, I think it should be 42 instead.
It takes T = 42/(3+4) = 6 hrs for them to meet.
And so Bob walked 6*4 = 24 miles.

Kenny